A dry cleaning business offers a pick-up and delivery service for a 10 percent surcharge. Management believes 60 percent of customers will take advantage of this service. They are also considering offering customers the option of opening an account and receiving monthly bills. They believe 60 percent of their customers ( regardless of whether or not they use the pick-up service ) will use the account service. If the two services are introduced to the market, what is the probability that a customer uses both services? ii.if the two services are introduced to the market, what is the probability that a customer uses only one of these services? iii.if the two services are introduced to the market, what is the probability that a customer uses neither of these services?






Answer by stanbon(75887)

A dry cleaning business offers a pick-up and delivery service for a 10 percent surcharge. Management believes 60 percent of customers will take advantage of this service. They are also considering offering customers the option of opening an account and receiving monthly bills. They believe 60 percent of their customers ( regardless of whether or not they use the pick-up service ) will use the account service.
------------------------------------
If the two services are introduced to the market, what is the probability that a customer uses both services?
-----
Ans: 0.6^2 = 0.36
------------------------------------
ii.if the two services are introduced to the market, what is the probability that a customer uses only one of these services?
Ans: 2*0.6*0.4 = 2*0.24 = 0.48
------------------------------------
iii.if the two services are introduced to the market, what is the probability that a customer uses neither of these services?
Ans: 1-(0.4)^2 = 1-0.16 = 0.84
--------
Cheers,
Stan H.
========